OpenStudy (anonymous):

For what values of x in [0,2π] does the graph of f(x)=x+2sinx have a horizontal tangent? List the values of x below.

OpenStudy (anonymous):

I don't know how to set up the equation, that's what I was asking.

OpenStudy (anonymous):

Find the first derivative, f'(x) Solve f'(x)=0 (horizontal tangent occur when the slope is zero). You can find x from here.

OpenStudy (swissgirl):

f'(x)=1+2Cos(x) 1+2Cos(x)=0 \(Cos(x)=\large\frac{-1} {2} \) \(Cos^{-1} \large(\frac{-1} {2})=x\) x=120 and 300

OpenStudy (swissgirl):

Did you follow?

OpenStudy (swissgirl):

I skipped out some steps but I am assuming you are familiar with trig

OpenStudy (swissgirl):

@Kalen Are u there?

OpenStudy (swissgirl):

Do u not follow?